OpenStudy (anonymous):

Use the graph to find the solution to the ions. Enter your answer as an ordered pair. y=-2x-4 y=x+2

OpenStudy (anonymous):

so for the second equation, it says y=x+2 we then plug x+2 in for y in the first equation so we then get x+2 = -2x-4 then we can solve for x and x =-3 THEN we plug that -3 back in for x to solve for y so y = (-3)+2 so y = -1 therefore the order pair will be (-3, -1)

OpenStudy (anonymous):

Thank u!